1]If A = B+C and the magnitude of A, B and C are 5,4,3 unit, the angle between A and B is :

(a)cos-13/5 (b) cos-14/5

(c)p/2 (d) sin-1 3/5

Grade:11

2 Answers

saumya shivhare
13 years ago

tan x= c sino/(b+ c coso)

on solving we get x= cos-1 4/5 or sin-1 3/5

vikas askiitian expert
13 years ago

A = B+C 

A-B = C , now squaring both sides

(A-B)2 = C2

 A2+B2-2(A.B) = C2

    (A.B) = (25+16-9)/2 = 32/2                                (A.B(A dot B) = ABCOS@)

    ABCOS@ = 16

       COS@ = 4/5

     SIN@ = 3/5 OR @ = SIN-3/5

OPTION D) IS CORRECT

approve my ans if u like it